OpenStudy (anonymous):

Integration by Parts ∫ tan^-1(y)dy

OpenStudy (anonymous):

\[\int\limits_{}^{} udv = uv - \int\limits_{}^{} vdu\] \[u = \tan^{-1}(y)\]\[du = \frac{ 1 }{ 1 + y^2 }\]\[dv = dy\]\[v = y\] plug and play. it will require a second, simpler integration

OpenStudy (anonymous):

small correction: \[du = \frac{ 1 }{ 1 + y^2 } dy\]***